The reaction between potassium chlorate and red phosphorus takes place when you strike a match on a matchbox. If you were to react 48.7 g of potassium chlorate (KClO3) with excess red phosphorus, what mass of tetraphosphorus decaoxide (P4O10) could be produced?
1
Expert's answer
2019-09-30T05:13:20-0400

The equation of chemical reaction (unbalanced!) is


"KCl{O_3} + P \to {P_4}{O_{10}} + KCl"

It's easy to balance it (first balance the "O" multiplying "KCl{O_3}" by 10 and "{P_4}{O_{10}}" by 3, then balance "K" multiplying "KCl" by 10 and balance "P" multiplying "P" by 12, "Cl" will be balanced automatically) and we get


"10KCl{O_3} + 12P \to 3{P_4}{O_{10}} + 10KCl"

Then calculate the molar mass of "KCl{O_3}" as the sum of atomic weights (can be found in periodic table) "{M_{KCl{O_3}}} = {M_K} + {M_{Cl}} + 3{M_O}" , we get


"{M_{KCl{O_3}}} \approx 39.098[\frac{{\text{g}}}{{{\text{mol}}}}] + 35.446[\frac{{\text{g}}}{{{\text{mol}}}}] + 3 \cdot 16[\frac{{\text{g}}}{{{\text{mol}}}}] = 122.544[\frac{{\text{g}}}{{{\text{mol}}}}]"

The same for "{P_4}{O_{10}}"


"{M_{{P_4}{O_{10}}}} \approx 4 \cdot 30.974[\frac{{\text{g}}}{{{\text{mol}}}}] + 10 \cdot 16[\frac{{\text{g}}}{{{\text{mol}}}}] = 283.896[\frac{{\text{g}}}{{{\text{mol}}}}]"

Now, we can see from the equation that for each 10 moles of "KCl{O_3}" we get 3 moles of "{P_4}{O_{10}}" , in other words for each 1 mole of "KCl{O_3}" we get "\frac{3}{{10}}" moles of "{P_4}{O_{10}}" . We can calculate the number of moles of "KCl{O_3}"


"{\nu _{KCl{O_3}}} = \frac{{48.7[{\text{g}}]}}{{122.544[\frac{{\text{g}}}{{{\text{mol}}}}]}} \approx 0.397[{\text{mol}}]"


And we can calculate the number of moles obtained "{P_4}{O_{10}}"


"{\nu _{{P_4}{O_{10}}}} = \frac{3}{{10}}{\nu _{KCl{O_3}}} = \frac{3}{{10}} \cdot 0.397[{\text{mol}}] \approx 0.12[{\text{mol}}]"

Let's calculate the mass of obtained "{P_4}{O_{10}}"


"{m_{{P_4}{O_{10}}}} = {\nu _{{P_4}{O_{10}}}} \cdot {M_{{P_4}{O_{10}}}} = 0.12[{\text{mol}}] \cdot 283.896[\frac{{\text{g}}}{{{\text{mol}}}}] \approx 34.07[{\text{g}}]"